# The chromatic number of finite projective spaces

### v1

@@ Abstract
The chromatic number of the finite projective space $\mathrm{PG}(n-1,q)$, denoted $\chi_{q}(n)$, is the minimum number of colors needed to color its points so that no line is monochromatic. We establish a recursive upper bound $\chi_{q}(n)\leq\chi_{q}(d)+\chi_{q}(n-d)$ for all $1\leq d<n$ and use it to prove new upper bounds on $\chi_{q}(n)$ for all $q$.
For $q=2$, we further refine this recursion and prove that
@@
We also consider $\chi_{q}(t;n)$, the minimum number of colors needed to color the points of $\mathrm{PG}(n-1,q)$ with no monochromatic (t1)(t-1)-dimensional subspace, and establish an equivalence between $\chi_{q}(t;n)$ and the multicolor vector-space Ramsey numbers $R_{q}(t;k)$. Using this equivalence together with our upper bounds on $\chi_{q}(t;n)$, we improve, for every fixed $t$, the best known lower bounds on $R_{q}(t;k)$ from $\Omega_{q}(\log k)$ to $\Omega(k)$. Finally, we improve the upper bound on $R_{2}(2;k)$ from a tower function of height $k$ to the function $k\log k$.
@@ 1 Introduction
It can be easily shown that $\chi_{2}(2)=2$. It is a standard exercise in the context of Property B and minimal non-2-colorable hypergraphs (see, for example, [31, Section 1.3]) that $\chi_{2}(3)=3$, that is, the chromatic number of the Fano plane is $3$. The next case is $\chi_{2}(4)=3$, which follows from the well-known partition of $\mathrm{PG}(3,2)$ into elliptic quadrics [9].
@@
Using our results, we thus obtain new lower bounds for $R_{q}(t;k)$ and also significantly improve the upper bound on $R_{2}(2;k)$.
@@
This improves the lower bounds on $R_{q}(t;k)$ from $\Omega(\log k)$ to $\Omega(k)$, and the upper bound on $R_{2}(2;k)$ from a tower function of height $k$ (see [11]) to the function $k\log k$. Note that for $t=2$, the lower bound $R_{q}(2;k)>k$ follows from $\chi_{q}(n)\leq n$, and thus the improved upper bounds in Theorem 3 give better lower bounds on $R_{q}(2;k)$. In particular, $R_{q}(2;k)>2k-O(1)$ for all fixed $q\geq 5$.
The paper is organized as follows. In Section 2 we introduce the general chromatic numbers $\chi_{q}(t;n)$, establish the basic recursion, and obtain upper bounds on $\chi_{q}(n)$. In Section 3 we focus on the binary case $q=2$, prove the recursion in Equation 1, and derive Theorem 1. In Section 4 we develop the connection to classical multicolor Ramsey numbers for triangles and determine exact values of $\chi_{2}(n)$ for small $n$. In Section 5 we make the link to multicolor vector space Ramsey numbers $R_{q}(t;k)$ and use our bounds on $\chi_{q}(t;n)$ to improve both lower and upper bounds for $R_{q}(t;k)$. In Section 6 we discuss future work and related open problems.
@@ 5 Connection to vector space Ramsey numbers
They further prove the following upper bounds for $R_{q}(t;k)$ when $q\in\{2,3\}$.
Theorem 6 ([11, Theorem 1.3]).
There exists a constant $C_{0}\approx 13.901$ such that for $\sigma_{2}=2$ and $\sigma_{3}=C_{0}$ the following holds. Let $q\in\{2,3\}$, let $k\geq 2$, and let $t_{k}\geq\cdots\geq t_{1}\geq 2$. Then $R_{q}(t_{1},\dots,t_{k})$ is at most a tower of height $\sum_{i=1}^{k-1}(t_{i}-1)+1$ of the form
$R_{q}(t_{1},\dots,t_{k})\leq\sigma_{q}^{\sigma_{q}^{\cdot^{\cdot^{\cdot^{\sigma_{q}^{3t_{k}}}}}}}.$
In particular, in the case $t_{1}=\dots=t_{k}=t$, the height of the tower is (k1)(t1)+1(k-1)(t-1)+1.
@@
We can also significantly improve the tower-type upper bound on this number for $q=t=2$.
@@ Acknowledgements
The authors thank Ferdinand Ihringer and Sam Mattheus for valuable insights and discussions.

### v2

@@ Abstract
The chromatic number of the finite projective space $\mathrm{PG}(n-1,q)$, denoted $\chi_{q}(n)$, is the minimum number of colors needed to color its points so that no line is monochromatic. We establish a recursive upper bound $\chi_{q}(n)\leq\chi_{q}(d)+\chi_{q}(n-d)$ for all $1\leq d<n$ and use it to prove new upper bounds on $\chi_{q}(n)$ for all $q$. For $q=2$, we further refine this recursion and prove that
@@
We also consider $\chi_{q}(t;n)$, the minimum number of colors needed to color the points of $\mathrm{PG}(n-1,q)$ with no monochromatic (t1)(t-1)-dimensional subspace, and establish an equivalence between $\chi_{q}(t;n)$ and the multicolor vector-space Ramsey numbers $R_{q}(t;k)$. Using this equivalence together with our upper bounds on $\chi_{q}(t;n)$, we improve, for every fixed $t$, the best known lower bounds on $R_{q}(t;k)$ from $\Omega_{q}(\log k)$ to $\Omega(k)$.
@@ 1 Introduction
It can be easily shown that $\chi_{2}(2)=2$. It is a standard exercise in the context of Property B and minimal non-2-colorable hypergraphs (see, for example, [32, Section 1.3]) that $\chi_{2}(3)=3$, that is, the chromatic number of the Fano plane is $3$. The next case is $\chi_{2}(4)=3$, which follows from the well-known partition of $\mathrm{PG}(3,2)$ into elliptic quadrics [9].
@@
Using our results, we thus obtain new lower bounds for $R_{q}(t;k)$.
@@
This improves the lower bounds on $R_{q}(t;k)$ from $\Omega(\log k)$ to $\Omega(k)$. Note that for $t=2$, the lower bound $R_{q}(2;k)>k$ follows from $\chi_{q}(n)\leq n$, and thus the improved upper bounds in Theorem 3 give better lower bounds on $R_{q}(2;k)$. In particular, $R_{q}(2;k)>2k-O(1)$ for all fixed $q\geq 5$.
@@
This improves the upper bound $R_{2}(2;k)=O(k\log^{7}k$) that follows from the more general bound given by Hunter and Pohoata [17, Theorem 1.4]), which was obtained using the breakthrough result of Kelley and Meka [18]. [^1] From the discussion around [17, Theorem 4.1] one can also deduce the upper bound of $O(k\log k)$ on $R_{2}(2;k)$ but it is not stated explicitly.
The paper is organized as follows. In Section 2 we introduce the general chromatic numbers $\chi_{q}(t;n)$, establish the basic recursion, and obtain upper bounds on $\chi_{q}(n)$. In Section 3 we focus on the binary case $q=2$, prove the recursion in Equation 1, and derive Theorem 1. In Section 4 we develop the connection to classical multicolor Ramsey numbers for triangles and determine exact values of $\chi_{2}(n)$ for small $n$. In Section 5 we establish the link to multicolor vector space Ramsey numbers $R_{q}(t;k)$ and use our bounds on $\chi_{q}(t;n)$ to improve the lower bounds for $R_{q}(t;k)$. In Section 6 we discuss future work and related open problems.
@@ 5 Connection to vector space Ramsey numbers
The best upper bounds on $R_{q}(t;k)$ in general are tower functions of height (k1)(t1)+1(k-1)(t-1)+1 [11, Theorem 1.3]. For $q=2$, Hunter and Pohoata [17] used the breakthrough result of Kelley and Meka [18] to show that $R_{2}(2;k)=O(k\log^{7}k)$. From the connection to $R(3;k)$ we immediately get the following.
@@ Acknowledgements
The authors thank Ferdinand Ihringer, Sam Mattheus, and Liana Yepremyan for valuable insights and discussions.